# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v3

@@ Abstract
[…]
@@ 1 Introduction
Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policy, leading to traffic challenges in passenger and cargo rail transport. The situation is aggravated by the fact that the rolling stock and especially railway infrastructure cannot keep up with the increase in transport needs, which makes railway systems overloaded.
Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays [23]. Examples of such disturbances include: late train departures/arrivals, extended dwell times, or (partial) track closures (also referred to as disruptions). These can last from several minutes up to hours. The impact of disturbances can propagate to multiple sections in the railway network (cf. the work of [50] and references therein). Thus, ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers, and train operators. To limit these impacts as much as possible it is necessary to make proper rescheduling decisions quickly. The dispatchers need to reschedule and partially reroute the trains, aiming at the minimization of negative consequences. Still, in many places, dispatchers are using their own intuition or simple heuristics like FCFS (First Come First Served); resulting in decisions far from the objective of consequence minimization. In the last decades, there has been a growing research interest in mathematical optimization methods in support of rail dispatchers in decision making. While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays [36], the maximal delay that cannot be avoided [14], or fuel consumption measures  [29].
As a railway network is a complex non-local structure, modeling a bigger portion of it is necessary for efficient suppression of the consequences of disturbances. Hence, large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost real time. Therefore it is vital to develop efficient algorithms. However, the current conventional rescheduling optimization models have difficulties addressing complex and large-scale instances in suitable computation time [8, 43, 35].
The railway dispatching/rescheduling problem can broadly be recognized as being equivalent to job-shop scheduling with blocking and no-wait constraints [49, 41]. A possible modeling approach is based on alternative graphs employing order and precedence variables [36], facilitating the formulation as a mixed integer linear program which is often large, hence, specialized algorithms are often used instead [35, 14]. An alternative approach is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants. Even though this results in very large problems, it is applied for both timetabling and dispatching/rescheduling [8, 39, 43, 18].
As time indexed models result in integer or 0-1 programs, they are suitable [53] for to be solved on new types of hardware: quantum annealers [2, 33]. These physical devices can be considered as stochastic heuristic solvers for Quadratic Unconstrained Binary Optimization (QUBO) problems, and there are few commercially available options, including D-Wave [31]. The exploration of their potential applications attracts a growing research interest (c.f. Section 2.2), in which railway applications are not yet strongly represented.
@@
In this paper, we address the train rescheduling problem in complex railway networks with mixed infrastructure including single, double, and multiple track railway lines with given planned train paths. Our consideration includes shunting movement of rolling stock between depots and stations followed by rolling stock connections. We apply a new hybrid rescheduling algorithm combining classical-quantum modelling and based on Quantum Annealing (QA). This work extends on a particular linear modeling strategy, partly explored on a toy model in [21]. We develop an Integer Linear Programming (ILP) model, which is solved with proprietary D-Wave solvers as well as with CPLEX for comparison. The D-Wave approach is analysed in detail, with respect to its applicability and performance in a practically relevant situation. The results suggest that quantum computing and QA in particular, although an early-stage technology, are ready for tackling challenging railway rescheduling problems. During the preparation of the present work, another contribution has appeared in the literature [56], which is similar to our in that it sets up a QUBO model for a somewhat similar problem, and solves it by a simulation of (quantum) annealing. This contribution, however, considers a different railway infrastructure: high-speed trains, and considers timetable optimization, for which it employs a model different from the one described here.
The main contributions of the present work are the following:
  • A railway traffic dispatching model is introduced for rescheduling trains due to disturbances and disruptions taking into account given timetable, rolling stock connections and a network with single or multi-track segments.
  • QA-based hybrid heuristics are, for the first time, applied to real-life problems of railway rescheduling optimization on urban railway networks, these heuristics have some but still meaningfully input of real QA.
  • The performance of the hybrid (quantum classical) D-Wave solver is demonstrated on a real-life network in Poland and diverse disturbance scenarios of different sizes and types.
  • The D-Wave hybrid solver provides good quality solutions in short time, also for the complex instances in limited computational time.
  • For certain instances the hybrid (quantum classical) D-Wave solver outperforms classical solvers in terms of computational time, yielding still feasible solutions.
The paper is organized as follows. In Section 2 we review state-of-the-art literature and identify scientific gaps we intend to fill, in Section 3 we describe problem under investigation, in Section 4 we evaluate our model, in Section 5 we discuss the hybrid solver we use, in Section 6 we present computational results.
2 Literature review
In this section, we present time-indexed models known from the literature, comparing them briefly with our modelling approach. For a general review of railway timetabling and rescheduling we refer to the reviews by [38] and [7]. Second, we review recent QA applications in optimization in general, and also in the rail/transport domain. Finally, we summarize the existing scientific gaps.
2.1 Classical railway dispatching/rescheduling
Time-indexed modelling of railway scheduling and rescheduling is quite common in literature for routing [39, 8] and scheduling [9, 29, 43, 13] . [8] focus on determining train’s path for a complex central railway station area (called the blocking - stairway), used a discrete-time model resulting in a $0-1$ program. Their model is successfully demonstrated on an operational day at the central railway station area Berne, Switzerland. Similarly, [39] consider train movements in a single major (and thus complex) railway junction, including a freight yard and a few minor stations. They build a mixed integer programming model that can also be solved close to optimal in case of practically relevant instances. They address the train movements on the level of detail of acceleration and deceleration strategies, hence adopting a significantly finer discrete time scale than our one-minute resolution. When compared to our problem, these two models addressed a high traffic in a complex station area. Instead, we are concerned more with a bigger area, multiple stations, and mixed track, but a lower train density and complexity. Our train routes fixed within the stations, and we also model shunting movements that is not a subject of the cited reference.
[9] proposed a graph theoretic formulation for the problem using a directed multigraph in which nodes correspond to departures/arrivals at a certain station at a given time instant. This formulation is used to derive an integer linear programming model that is solved using a Lagrangian relaxation. [13] introduced a new pure 0-1 programming formulation, and call it Tick Formulation, to model the Deadlock Detection (DD) problem. [43] focused “on the train dispatching problem on an N-track network, with the main challenge of how to formulate specific retiming, reordering, retracking and rerouting options in combination”. Their model was demonstrated on generic networks of different structure than what we address. In our situation we are limited to retiming and reordering of trains. Structurally, their model shows similarities to ours, including the role of binary variables and also the introduction of "cells" that generalize the objects the occupancy of which can be controlled: they can be one or more blocks, stations, etc. In our model, the decision variables will be explicitly linked to a subset of stations.
2.2 Quantum Annealing and its applications
Quantum Annealing (QA) [15] is an optimization method analogous to Simulated Annealing (SA). Both can be used for the unconstrained minimization of a quadratic function of ±$\pm 1$-valued variables, which represents the energy configuration for a set of spins of the Ising model [30, 4]; an important and celebrated problem in physics. The Ising problem can be equivalently rewritten in a form of a Quadratic Unconstrained Binary Optimization (QUBO) problem with binary ($0-1$) variables. The Max-Cut problem [5] is also equivalent to the Ising model or QUBO. The problem is NP-hard, has tremendous literature in classical optimization, and there is significant ongoing progress in the development of algorithms to solve it [22, 28]. An extensive list of Ising problems and their formulations are presented e.g. in the work of [37].
Hardware quantum annealers, like the D-Wave machine, implement a quantum version of the Ising model, assigning a quantum bit, i.e. a two-level physical quantum system to each bit. The system employs real physical two-level quantum systems with a tunable energy operator (Hamiltonian), the energy being the objective function of the optimization problem. An adiabatic evolution of the quantum system is implemented: the interactions between the spins are slowly changed from an initial Hamiltonian with a simple minimal-energy state to the Hamiltonian corresponding to the objective according to the rules of quantum mechanics. According to the adiabatic theorem of quantum mechanics [3], under certain conditions the physical system remains in its lowest energy state during the evolution, and an optimal, minimum-energy solution can be possibly read out at the end.
An ideal adiabatic quantum computer would be a device with a large number of perfect quantum bits completely decoupled from their environment, very close to zero temperature. Any pair of qubits could be coupled, and even, more general couplings, e.g. involving $3$ or more spins could take place. Such a system would exhibit perfect quantum coherence and entanglement. Even in such an ideal system the evolution time to reach at a minimum energy state with high enough probability is inverse proportional to the gap between the minimal energy configuration and the one closest to it. Certainly the gap is not known in advance, hence, the annealing time has to be determined by experimenting in practice. A more severe issue is that the gap can be very small , hence, it is not possible to solve all possible hard problems on such a setup. In spite of that, there are problems in which this approach can be efficient.
[…]
Operationally, from the optimizer’s point of view quantum annealers can be viewed as probabilistic heuristic solvers returning a statistical sample [20] of configurations which are supposed to be optimal or close to optimal. It must be stressed that the quantum annealers are not algorithms running on digital computers; they are analog devices implemented physically. This latter implies that the coefficients of the problem are encoded with a limited accuracy, and the algorithmic properties of the particular optimization problem such as its complexity class will not determine the solver’s behaviour.
To overcome the problem of limited size and accuracy, quantum annealers of the present state of the art are often used in hybrid (quantum classical) solvers, orchestrating classical algorithms and using QA as a subroutine in order to address hard problem instances more efficiently. Solvers available in D-Wave’s ’Leap Hybrid Solver Service (HSS)’ [11], including the one used in the present work, belong to this family. Meanwhile, quantum technology keeps on developing, systems of bigger size and better topology are regularly announced, they are more and more affordable, and there is a growing community around them.
Currently, the applicability of QA technologies is being explored. Benchmark problems are solved [55, 42], comparisons with digital computers are made [32], etc. The application of quantum annealers in transportation optimization is a new area with only a few contributions so far. There is an apparent interest in this research direction in the aviation domain [47, 51, 48], which relates also to the already mentioned benchmarking of annealers as done by[55]. Shipment rerouting was also addressed in the context of QA [58]. There are a few examples of hybrid solvers’ applications, too. These include traffic flow optimization [44], multi-car paint shop optimization [57], tail assignment problem [40], and vehicle optimization [24].
In railway context, the first applications of QA can be found in train dispatching/rescheduling [19, 21] and rolling stock planning [27]. In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us [18]. A followup paper [21] laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems. The present work continues this line of research.
2.3 Scientific gaps
@@
We recognize several important scientific gaps. First, existing optimization models suffer from the curse of dimensionality, and inability to solve larger real-life instances. Second, the QA models for the studied optmimzation problem that have been introduced so far [19, 21] had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices. Third, no hybrid QA-based models have been used for real-life railway, or even other schedule-based modes, like public transport and air traffic planning and/or rescheduling.
In this paper we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of a urban railway networks. To do so, we use a hybrid approach combining quantum and classical computing. For a fair comparison of current and future QA approaches with state-of-the-art classical approaches, we elaborate the railway model readily suitable for both classical and hybrid (quantum classical) solvers in a comparable manner. Quantum computing is believed to develop rapidly in near future. Hence, demonstration quantum readiness for railway problems right now is step forward towards the demonstration of quantum advantage in future. On future quantum devices the comparison with classical solvers is expected to be in strong favour for the quantum ones. In this way we also contribute in developing a novel general set of railway rescheduling models that has a potential to scale well with the size of the problem, aiming to overcome the curse of dimensionality.
@@ 3 Problem description
We consider a railway rescheduling problem that includes train reordering, retiming, and shunting movements with rolling stock circulation at stations. Also, we consider an urban railway network with mixed tracks from single, double, up to quadruple tracks.
We model a railway network with edges and nodes of a graph. Figure 1 depicts an exemplary network layout composed of nodes (station or junction) and edges (single, double, multiple track lines). Each edge is composed of one or more tracks. Each track consists of blocks sections defined by pairs of signals [39]. Each block section can be occupied by at most one train at a time. A subsequent train is allowed to enter the block only after a minimal headway: minimal time span between the trains. We assume a 2-block signalling system, meaning that two free blocks are required between the consecutive trains. We also assume green way policy (trains have the free way to move at the maximal allowed speed between stations), resulting in constant running times for each block.
Each node is composed of station tracks (blocks) and interlocking areas. One station track can be occupied by one train at a time. Routing dependencies between pairs of trains competing for the same resource in interlocking areas are considered in order to to guarantee that only one train from the pair can occupy the area at the time.
A train’s route is the sequence of blocks the train passes during its journey. A train path is a sequence of arrival and departure times of a particular train assigned to a train route. We assume a one-minute resolution for all time parameters such as timetable time, running and dwell time minimal headway time or passing time, resulting in integer variables. This assumption of discretized time is needed to enable the use of quantum-based solvers, and, more importantly, introduces the possibility of pruning the inherently binary variables. The relaxation of integer constraints on time variables is not expected to improve computation time significantly, as the real difficulty is tied to precedence of trains, encoded later on precedence binary variables.
In the network, two trains can follow each other, i.e. keep the given order, meet and pass (M-P) when going the opposite directions, and meet and overtake (M-O), i.e. change the order, when going the same direction. Single tracks are designed for bidirectional traffic, double-tracks for unidirectional, one for each direction, and multi-track lines can be combination of unidirectional and bidirectional. On single-track lines, M-Ps and M-Os are only possible at stations. To prevent M-P on single-track lines, we determine the set $\mathcal{J}^{2}_{\text{single}}$ as the set of all pairs of trains that can potentially meet on the single track line heading in opposite directions.
@@
The train traffic is scheduled on the basis of the given timetable. The timetable contains all the train paths. In the train rescheduling problem, we are given an initial timetable that is conflict-free. However, conflicts may appear due to disturbances such as late departures and/or arrivals due to excessive passenger demand, malfunctioning rolling stock, etc. Following [10], a conflict is an inadmissible situation when at least a pair of trains claim the same resource (e.g., block section, switch) simultaneously. We assume that the conflict may occur either on the railway line or at the station. Possible conflicts that can occur on the line include the lack of the minimum headway between two subsequent trains heading on the same track in the same direction, or claiming a segment of a single track line by two trains heading in opposite directions at the same time. Conflicts at stations include claiming a station track by two trains at the same time, claiming a station switch (in the interlocking area) by two trains at the same time.
@@
The conflicts have to be solved by modifying the original timetable, applying decisions on the train sequencing and retiming for trains claiming for the same resource. Such an intervention in the structure of train paths implies additional changes to maintain the feasibility of the modified timetable. These changes typically result in delays of additional trains that were not directly involved in the conflict otherwise, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable. It is common in the literature to minimize a function of these secondary delays (see e.g. the work of [14]); we will do this in the present contriubution, too).
@@
An important element of our approach is that from amongst all stations we first determine a subset of the decision stations; we assume that the direct decisions implied by our model affect these stations only. Hence, as decision stations we select those stations, where routes of trains intersect, where trains start or terminate, or where selected part network is bounded. The motivation is that if the routes of trains are fixed, the decisions on modifying their train paths has to be made with respect to these stations. (If we change some of the trains’ routes in a re-routing process, new model with new parameters if developed, e.g. additional decision stations may appear.) In what follows, by a station we always mean a decision station. This also means that non-decision stations appear only through the parameter values in the model, they do not appear as indices of decision variables or parameters. Headways, for instance, are calculated between decision stations, taking into account all the line blocks and station blocks of non-decision stations in between.
@@
On the station where a train terminates or sets off, shunting is also modeled. The goal of shunting is to move the train from the passengers’ service track to the depot or vice versa. The depot is treated as the station, and consider it as a black box, without detailed layout. We treat shunting movements as service train from depot to the starting station of the service train, or from the terminating station to the depot. The rolling stock circulation condition is applied to ensure the precedence between the service train and the actual train.

### v4

@@ Abstract
railway rescheduling, conflict management, heterogeneous urban railway network, quantum annealing, hybrid quantum-classical heuristics
@@ 1 Introduction
Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policy, leading to traffic challenges in passenger and cargo rail transport. The situation is aggravated by the fact that the rolling stock and especially railway infrastructure cannot keep up with the increase in transport needs, which overloads railway systems. Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances and disruptions in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays (Ge et al., 2022). Examples of such disturbances include: late train departures/arrivals, extended dwell times. These can last from several minutes up to hours. A disturbance is a smaller perturbation that can be handled solely by modifying the existing train paths, whereas a disruption affects rolling stock and crew schedules, too (Cacchiani et al., 2014). In the present work we focus on rescheduling of trains, which is the network operator’s perspective, even though we also consider (partial) track closures, which are are typically considered as disruptions. In our examples, however they will be managed by modifying existing train paths. The impact of disturbances and disruptions can propagate to multiple sections in the railway network (c.f. Törnquist (2007) and references therein). Thus, ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers, and train operators.
As a railway network is a complex non-local structure, modeling a bigger portion of it is necessary for efficient suppression of the consequences of disturbances. While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays (Lange and Werner, 2018), the maximal delay that cannot be avoided (D’Ariano et al., 2007), or fuel consumption measures  (Harrod, 2011). Large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost real time. The railway dispatching/rescheduling problem is recognized as equivalent to job-shop scheduling with blocking and no-wait constraints (Szpigel, 1973; Mascis and Pacciarelli, 2002). It has an extensive literature, already summarized in a number of reviews, including those by Cacchiani et al. (2014), by Lamorgese et al. (2018), or by Philip and Swapnesh (2022).
The maybe most frequently applied modeling strategy is based on alternative graphs (D’Ariano et al., 2007; Mascis and Pacciarelli, 2002) which can be formulated as as a mixed integer linear program, often referred to as the big-M model. As this model often becomes too big to be solved on time, a number of strategies are applied, including heuristics, specific branch-and-bound methods, alternative formulations, decompositions, and the combinations of these; see the work of Lamorgese et al. (2018) for a review of the model and the methods.
An alternative approach is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants. Even though this results in very large problems, it is applied for both timetabling and dispatching/rescheduling (Caimi et al., 2012; Lusby et al., 2013; Meng and Zhou, 2014). As this approach results in a $0-1$ program, which is promising when one considers the use of $0-1$ solvers such as quantum hardware. Indeed, job-shop problems addressed on pure quantum hardware (Venturelli et al., 2015), and even in certain hybrid approaches (Kurowski et al., 2020) use similar models, and our previous work (Domino et al., 2023) which was the first to apply a quantum solver to railway dispatching problems also followed this route. In the present work, however, we deal with a variant of the big-M model, albeit with discrete time variables, resulting in an integer linear program (ILP). We introduce a variant of this type of model which maintains a reasonable model size while addressing practically relevant problem instances. In this way, a hybrid classical-quantum integer solver can be applied to solve them, and a comparison with a classical solver is also possible.
@@
In particular, we address the train rescheduling problem in complex railway networks with mixed infrastructure including single, double, and multiple-track railway lines with given planned train paths. Our consideration includes shunting movement of rolling stock between depots and stations followed by rolling stock connections. We apply a new hybrid rescheduling algorithm combining classical-quantum modeling and based on Quantum Annealing (QA). This work extends on a particular linear modeling strategy, partly explored on a toy model by Domino et al. (2022b). Our (ILP) model, tailored for the particular practical railway situation, is solved with proprietary D-Wave solvers as well as with CPLEX for comparison.
@@
We recognize several important scientific gaps. First, existing optimization models suffer from the curse of dimensionality, and inability to solve larger real-life instances. Second, the QA models for the studied optimization problem that have been introduced so far had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices. Third, no hybrid QA-based models have been used for real-life railway, or even other schedule-based modes, like public transport and air traffic planning and/or rescheduling. In this paper we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of a urban railway networks.
The paper is organized as follows. In Section 2 we briefly review state-of-the-art of quantum annealing and its applications. In Section 3 we describe problem under investigation, in Section 4 we evaluate our model, in Section 5 we discuss the hybrid solver we use, in Section 6 we present computational results.
2 Quantum annealing
Quantum computing devices became available for practical computational purposes in the last few years. Quantum annealers are such devices: they implement two-level quantum systems (quantum bits) with tunable pairwise interactions: an Ising system well-known in physics (Ising, 1925; Bian et al., 2010). The expression of the energy of such a system is one of the many equivalent forms of quadratic unconstrained binary optimization problems (QUBOs): the extremum of a quadratic function of binary variables, which is an NP-hard problem in general.
Hence, a QUBO objective can be expressed as the energy of the Ising system: the configuration of the pairwise couplings of the spins and the applied local fields, which are tunable in quantum hardware. According to the adiabatic theorem of quantum mechanics (Avron and Elgart, 1999), tuning the system’s configuration from a suitably chosen initial state to the one corresponding to the actual objective slowly enough, under certain conditions the optimal configuration can be read out. Unfortunately, not all QUBOs meet the required conditions, so even in an ideal setting the quantum annealer is not a universal solver of hard problems, but it can be efficient in certain problems. Quantum annealing can be also viewed as a process analogous to simulated annealing, in which the noise required to avoid staying in local optima has quantum (as opposed to thermal) origins  (Das and Chakrabarti, 2008).
In real physical hardware, like the D-Wave machine, thermal noise and imperfections cannot be avoided, hence, the physical devices realize a combination of thermal and quantum annealing, which renders them a probabilistic heuristic solver for QUBOs. They output a number of configurations typically close to, and optionally including the optimal one.
The class of QUBO problems is well studied and have a number of applications; we refer the reader to a recent monograph edited by Punnen (2022). Linear unconstrained binary problems can also be considered as a special case of QUBOs, while linear and quadratic constraints can be taken into account with penalties (Gusmeroli et al., 2022). Using appropriate binary encoding of the variables, integer programming problems can also be treated. Hence, an efficient QUBO solver can be useful in many problems, including the integer linear programs arising in the present contribution.
The state-of-the-art physical quantum annealers are smaller systems of a few hundred quantum bits and not all pairs are coupled. The fixed topology of qubit couplings means that the problem’s graph defined by the nonzero coupling has to be embedded as an induced subgraph of the topology of the system. This embedding is a hard problem itself (Zbinden et al., 2020). It often requires to couple multiple physical qubits to represent a logical bit of the problem. As quantum annealers are not algorithms running on digital computers but analog devices implemented physically, the coefficients of the problem are encoded with a limited accuracy.
To overcome the problem of limited size and accuracy, quantum annealers of the present state of the art are often used in hybrid (quantum-classical) solvers, orchestrating classical algorithms and using QA as a subroutine in order to address hard problem instances more efficiently. Solvers available in D-Wave’s ’Leap Hybrid Solver Service (HSS)’ (D-Wave Quantum Inc., 2022a), including the one used in the present work, belong to this family. Meanwhile, quantum technology keeps on developing, systems of bigger size and better topology are regularly announced, they are more and more affordable, and there is a growing community around them.
Currently, the applicability of QA technologies is being explored. In the railway context, the first applications of QA can be found in train dispatching/rescheduling (Domino et al., 2022b) and rolling stock planning (Bickert et al., 2021). In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us (Domino et al., 2023). A follow-up paper (Domino et al., 2022b) laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems. The present work continues this line of research, which is gaining increasing attention, e.g. the recent work of Xu et al. (2023).
@@ 3 Problem description
We consider a railway rescheduling problem that includes train reordering, retiming, and shunting movements with rolling stock circulation at stations. Also, we consider an urban railway network with mixed tracks from single, double, to quadruple tracks.
We model a railway network with edges and nodes of a graph. Figure 1 depicts an exemplary network layout composed of nodes (station or junction) and edges (single, double, multiple-track lines). Each edge is composed of one or more tracks. Each track consists of block sections defined by pairs of signals (Lusby et al., 2013). Each block section can be occupied by at most one train at a time. A subsequent train is allowed to enter the block only after a minimal headway: minimal time span between the trains. We assume a 2-block signaling system, meaning that two free blocks are required between the consecutive trains. We also assume a green way policy (trains have the free way to move at the maximal allowed speed between stations), resulting in constant running times for each block.
Each node is composed of station tracks (blocks) and interlocking areas. One station track can be occupied by one train at a time. Routing dependencies between pairs of trains competing for the same resource in interlocking areas are considered to guarantee that only one train from the pair can occupy the area at the time.
A train’s route is the sequence of blocks the train passes during its journey. A train path is a sequence of arrival and departure times of a particular train assigned to a train route. We assume a one-minute resolution for all time parameters such as timetable time, running and dwell time minimal headway time, or passing time, resulting in integer variables. This assumption of discretized time is needed to enable the use of quantum-based solvers, and, more importantly, introduces the possibility of pruning the inherently binary variables. The relaxation of integer constraints on time variables is not expected to improve computation time significantly, as the real difficulty is tied to the precedence of trains, encoded later on precedence binary variables.
In the network, two trains can follow each other, i.e. keep the given order, meet and pass (M-P) when going the opposite directions, and meet and overtake (M-O), i.e. change the order, when going the same direction. Single-track lines are designed for bidirectional traffic, double track lines for unidirectional, one for each direction, and multi-track lines can be a combination of unidirectional and bidirectional. On single-track lines, M-Ps and M-Os are only possible at stations. To prevent M-P on single-track lines, we determine the set $\mathcal{J}^{2}_{\text{single}}$ as the set of all pairs of trains that can potentially meet on the single-track line heading in opposite directions.
@@
The train traffic is scheduled based on the given timetable. The timetable contains all the train paths. In the train rescheduling problem, we are given an initial timetable that is conflict-free. However, conflicts may appear due to disturbances such as late departures and/or arrivals due to excessive passenger demand, malfunctioning rolling stock, etc. Following Corman et al. (2012), a conflict is an inadmissible situation when at least a pair of trains claim the same resource (e.g., block section, switch) simultaneously. We assume that the conflict may occur either on the railway line or at the station. Possible conflicts that can occur on the line include the lack of the minimum headway between two subsequent trains heading on the same track in the same direction, or claiming a segment of a single-track line by two trains heading in opposite directions at the same time. Conflicts at stations include claiming a station track by two trains at the same time, or claiming a station switch (in the interlocking area) by two trains at the same time.
@@
The conflicts have to be solved by modifying the original timetable, applying decisions on the train sequencing, and retiming for trains claiming the same resource. Such an intervention in the structure of train paths implies additional changes to maintain the feasibility of the modified timetable. These changes typically result in delays of additional trains that were not directly involved in the conflict otherwise, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable. We will minimize a function of these secondary delays.
@@
An important element of our approach is that from among all stations, we first determine a subset of the decision stations; we assume that the direct decisions implied by our model affect these stations only. Hence, as decision stations, we select those stations, where routes of trains intersect, where trains start or terminate, or where the selected part of the network is bounded. The motivation is that if the routes of trains are fixed, the decisions on modifying their train paths have to be made with respect to these stations. (If we change some of the trains’ routes in a re-routing process, the new model with new parameters is developed, e.g. additional decision stations may appear.) In what follows, by a station we always mean a decision station. This also means that non-decision stations appear only through the parameter values in the model, they do not appear as indices of decision variables or parameters. Headways, for instance, are calculated between decision stations, taking into account all the line blocks and station blocks of non-decision stations in between.
@@
On the station where a train terminates or sets off, shunting is also modeled. The goal of shunting is to move the train from the passengers’ service track to the depot or vice versa. The depot is treated as the station, and consider it as a black box, without a detailed layout. We treat shunting movements as a service train from the depot to the starting station of the service train, or from the terminating station to the depot. The rolling stock circulation condition is applied to ensure the precedence between the service train and the actual train.